# Weak-to-Strong Search: Align Large Language Models via Searching over Small Language Models

### v1

@@ 6 Discussion
Given language models’ autoregressive nature, we can also view language model alignment as solving a token-level MDP [20]. This token-level MDP is define by the tuple $(\mathcal{S},\mathcal{A},f,r({\mathbf{s}}_{t},{\mathbf{a}}_{t}))$. Here, the state ${\mathbf{s}}_{t}:=({\mathbf{x}},{\mathbf{y}}_{<t})\in\mathcal{S}$ consists of the prompt and all response tokens generated so far; the action ${\mathbf{a}}:=y_{t}$ determines the next token to generate from the vocabulary $\mathcal{A}$; the dynamics $f$ is a deterministic function that updates the state by concatenating the current state and action ${\mathbf{s}}_{t+1}:=({\mathbf{s}}_{t},{\mathbf{a}}_{t})$; $r({\mathbf{s}}_{t},{\mathbf{a}}_{t})$ is a sparse reward that equals $r({\mathbf{x}},{\mathbf{y}})$ if ${\mathbf{a}}_{t}$ is EOS and 00 otherwise. We use $\rho_{\pi}({\mathbf{s}}_{t})$ to denote the state marginals of the trajectory distribution induced by the policy $\pi$. Under the token-level MDP, the objective from Eq. 7 can be written as
@@
$\sum_{t=1}^{H}\beta\log\frac{\pi^{*}({\mathbf{a}}_{t}\mid{\mathbf{s}}_{t})}{\pi_{\text{ref}}({\mathbf{a}}_{t}\mid{\mathbf{s}}_{t})}=\begin{cases}-V^{*}({\mathbf{s}}_{1})+V^{*}(({\mathbf{s}}_{H},{\mathbf{a}}_{H})),&\text{ if }{\mathbf{a}}_{H}\text{ is not {EOS}}\\ -V^{*}({\mathbf{s}}_{1})+r({\mathbf{s}}_{H},{\mathbf{a}}_{H}),&\text{ if }{\mathbf{a}}_{H}\text{ is {EOS}}\\ \end{cases}$ (13)
where $V^{*}({\mathbf{s}}_{H+1})=V^{*}(({\mathbf{s}}_{H},{\mathbf{a}}_{H}))$ due to the deterministic transition. Eventually, in the context of sequence-level MDP (Section 4), where we define ${\mathbf{y}}$ as a complete response, ${\mathbf{y}}^{\prime}$ as a response that can be either complete or incomplete, we have that ${\mathbf{s}}_{1}=({\mathbf{x}},{\mathbf{y}}_{<1})=({\mathbf{x}},\varnothing)={\mathbf{x}}$ and $({\mathbf{s}}_{H},{\mathbf{a}}_{H})=(({\mathbf{x}},{\mathbf{y}}_{<H}),{\mathbf{y}}_{H})=({\mathbf{x}},{\mathbf{y}}_{<H+1})$. Thus, we can rewrite Eq. 13, with a slight abuse of notations, as
$\log\frac{\pi^{*}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})}\propto\begin{cases}-V^{*}({\mathbf{x}})+V^{*}({\mathbf{x}},{\mathbf{y}}^{\prime}),&\text{ if }y^{\prime}_{|{\mathbf{y}}^{\prime}|}\neq\texttt{EOS}\text{ (${\mathbf{y}}^{\prime}$ is incomplete)}\\ -V^{*}({\mathbf{x}})+r({\mathbf{x}},{\mathbf{y}}^{\prime}),&\text{ if }y^{\prime}_{|{\mathbf{y}}^{\prime}|}=\texttt{EOS}\text{ (${\mathbf{y}}^{\prime}$ is complete)}\end{cases}$ (14)
Appendix B Changelog
[…]
Appendix C Further Details on the Experimental Setup
C.1 Controlled-Sentiment Generation & Summarization
C.1.1 Model Specification
@@ C.1.2 Hyperparameters Specification
C.1.3 Compute Resources Specification
@@
C.1.4 Gold Reward Model Details
@@
For controlled-sentiment generation, we reuse the publicly available distilbert-imdb as the gold reward model $r_{\text{gold}}$, which is a fine-tuned distilbert-base-uncased [52] on the imdb dataset [53] for classifying movie review sentiments. We define the gold reward $r_{\text{gold}}$ as $\log p(\text{positive}\,|\,x,y)-\log p(\text{negative}\,|\,x,y)$ to encourage positive review. Synthetic preferences are collected using the truncated movie reviews as prompts ${\mathbf{x}}$, and pairwise completions from gpt2-imdb, ranked with $p({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}\mid{\mathbf{x}})=\sigma(r_{\text{gold}}({\mathbf{x}},{\mathbf{y}}_{1})-r_{\text{gold}}({\mathbf{x}},{\mathbf{y}}_{2}))$, as preferences.
@@
C.1.5 Direct Tuning Details
@@
C.1.6 Prompt Template for Sampling from Base Models
@@
C.2 Instruction Following
C.2.1 Model Specification
@@
C.2.3 Compute Resources Specification.
@@
Appendix D Extended Experimental Results
D.1 Chunk-level Beam Search Ablations
@@
Figure 7: L ablations for CBS (W, K = 4, 4). We show the mean rewards ($\pm$ 2 standard deviations).
@@
Figure 8: W, K ablations for CBS ($\mathbf{L=5}$). We show the mean rewards across three random seeds. With the same computation budget (i.e., same $W\cdot K$) the optimal hyperparameters differ by tasks.
D.2 Evaluation Results for Instruction Following
@@
Appendix E Sample Generations
E.1 Controlled-Sentiment Generation Sample Generations
@@
E.2 Summarization Sample Generations

### v3

@@ Acknowledgements
Given the autoregressive nature of language models, we can frame language model alignment as solving a token-level MDP [20]. This token-level MDP is define by the tuple $(\mathcal{S},\mathcal{A},f,r({\mathbf{s}}_{t},{\mathbf{a}}_{t}))$. Here, the state ${\mathbf{s}}_{t}:=({\mathbf{x}},{\mathbf{y}}_{<t})\in\mathcal{S}$ consists of the prompt and all response tokens generated so far; the action ${\mathbf{a}}:=y_{t}$ determines the next token to generate from the vocabulary $\mathcal{A}$; the dynamics $f$ is a deterministic function that updates the state by concatenating the current state and action ${\mathbf{s}}_{t+1}:=({\mathbf{s}}_{t},{\mathbf{a}}_{t})$; $r({\mathbf{s}}_{t},{\mathbf{a}}_{t})$ is a sparse reward that equals $r({\mathbf{x}},{\mathbf{y}})$ if ${\mathbf{a}}_{t}$ is EOS and 00 otherwise. We use $\rho_{\pi}({\mathbf{s}}_{t})$ to denote the state marginals of the trajectory distribution induced by the policy $\pi$. Under the token-level MDP, the objective from Eq. 7 can be written as
@@
$\sum_{t=1}^{H}\beta\log\frac{\pi^{*}({\mathbf{a}}_{t}\mid{\mathbf{s}}_{t})}{\pi_{\text{ref}}({\mathbf{a}}_{t}\mid{\mathbf{s}}_{t})}=\begin{cases}-V^{*}({\mathbf{s}}_{1})+V^{*}(({\mathbf{s}}_{H},{\mathbf{a}}_{H}))&\text{ if }{\mathbf{a}}_{H}\text{ is not {EOS}}\\ -V^{*}({\mathbf{s}}_{1})+r({\mathbf{s}}_{H},{\mathbf{a}}_{H})&\text{ if }{\mathbf{a}}_{H}\text{ is {EOS}},\\ \end{cases}$ (13)
where $V^{*}({\mathbf{s}}_{H+1})=V^{*}(({\mathbf{s}}_{H},{\mathbf{a}}_{H}))$ due to the deterministic transition. Now, returning back to the sequence-level MDP (Section 4), where we define ${\mathbf{y}}$ as a complete response, ${\mathbf{y}}^{\prime}$ as a response that can be either complete or incomplete, we have that ${\mathbf{s}}_{1}=({\mathbf{x}},{\mathbf{y}}_{<1})=({\mathbf{x}},\varnothing)={\mathbf{x}}$ and $({\mathbf{s}}_{H},{\mathbf{a}}_{H})=(({\mathbf{x}},{\mathbf{y}}_{<H}),{\mathbf{y}}_{H})=({\mathbf{x}},{\mathbf{y}}_{<H+1})$. Thus, we can rewrite Eq. 13, with a slight abuse of notations, as
$\log\frac{\pi^{*}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})}\propto\begin{cases}-V^{*}({\mathbf{x}})+V^{*}({\mathbf{x}},{\mathbf{y}}^{\prime})&\text{ if }{\textnormal{y}}^{\prime}_{|{\mathbf{y}}^{\prime}|}\neq\texttt{EOS}\text{ (${\mathbf{y}}^{\prime}$ is incomplete)}\\ -V^{*}({\mathbf{x}})+r({\mathbf{x}},{\mathbf{y}}^{\prime})&\text{ if }{\textnormal{y}}^{\prime}_{|{\mathbf{y}}^{\prime}|}=\texttt{EOS}\text{ (${\mathbf{y}}^{\prime}$ is complete)}.\end{cases}$ (14)
Appendix B Further Details on the Experimental Setup
B.1 Controlled-Sentiment Generation & Summarization
B.1.1 Model Specification
@@ B.1.2 Hyperparameters Specification
B.1.3 Compute Resources Specification
@@
B.1.4 Gold Reward Model Details
@@
For controlled-sentiment generation, we reuse the publicly available distilbert-imdb to define the gold reward model $r_{\text{gold}}$. Distilbert-imdb is a fine-tuned classifier $p$ on the imdb dataset [53] to classify movie review sentiments. We define the gold reward $r_{\text{gold}}$ as $\log p(\text{positive}\,|\,x,y)-\log p(\text{negative}\,|\,x,y)$ to encourage positive review. Synthetic preferences are collected using the truncated movie reviews as prompts ${\mathbf{x}}$, and pairwise completions from gpt2-imdb, ranked with $p({\mathbf{y}}_{1}\succ{\mathbf{y}}_{2}\mid{\mathbf{x}})=\sigma(r_{\text{gold}}({\mathbf{x}},{\mathbf{y}}_{1})-r_{\text{gold}}({\mathbf{x}},{\mathbf{y}}_{2}))$, as preferences.
@@
B.1.5 Direct Tuning Details
@@
B.1.6 Prompt Template for Sampling from Base Models
@@
B.2 Instruction Following
B.2.1 Model Specification
@@ B.2.2 Hyperparameters Specification
B.2.3 Compute Resources Specification
@@
Appendix C Extended Experimental Results
C.1 Chunk-level Beam Search Ablations
@@
Figure 7: L ablations for CBS (W, K = 4, 4). We show the mean rewards ($\pm$ standard deviations).
@@
Figure 8: W, K ablations for CBS ($\mathbf{L=5}$). We show the mean rewards across three random seeds. With the same computation budget (i.e., same $WK$), the optimal hyperparameters differ by tasks.
C.2 Evaluation Results for Instruction Following
@@
C.3 RL Fine-tuning with Language Model Reward
In Section 5, we have demonstrated that by converting weak language models to a per-token dense reward function (Eq. 4), we can further improve strong models at test time without additional training. This section presents preliminary experiments showing that this reparametrized dense reward function can also benefit RL fine-tuning. Using the same setup from Section 5, we start with a dense reward function parametrized by tuned and unturned gpt2 pair $(\pi^{*},\pi_{\text{ref}})$ and then we tune two larger models under this dense reward on sentiment-controlled generation using PPO [57]. Figure 10 shows the fine-tuning results with different reward sparsity, where we find that dense reward function does benefit RL fine-tuning, and weak-to-strong generalization [21] also occurs consistently during training when the reward signal comes from weaker language models (note that $\pi^{*}$ only achieve a $4.6$ gold reward on sentiment-controlled generation from Figure 3).
(a) PPO fine-tuning results for gpt2-large.
(b) PPO fine-tuning results for Llama-2-7b w/ LoRA ($r=128$, $\alpha=128$).
Figure 10: PPO fine-tuning with chunk-level dense rewards for controlled-sentiment generation. Using dense reward parameterized by tuned and untuned gpt2 models (see Section 5 for details), we train larger base models with PPO. The chunk length $L$ controls the reward sparsity. For example, $L=5$ means rewards are accumulated and emitted every 5 tokens (delayed to the last token of each chunk) while $L=\text{inf}$ corresponds to vanilla PPO with sequence-level sparse rewards. Denser rewards facilitate credit assignment and accelerate training, improving both the achieved return on training prompts (cumulative dense rewards over the complete responses) (left) and the achieved gold reward on validation prompts (right). We plot mean $\pm$ std over three random seeds.
Appendix D Sample Generations
D.1 Controlled-Sentiment Generation Sample Generations
@@
D.2 Summarization Sample Generations